OpenStudy (anonymous):

How can 1/6x − 5 = 1/5x + 2 be set up as a system of equations?

OpenStudy (astrophysics):

Are you solving for x?

OpenStudy (astrophysics):

You just gave 1 equation

OpenStudy (anonymous):

@Astrophysics 1. 6 y − x = −30 5y − x = 10 2. 5y + x = −30 5y + x = 10 3. 6y − 6x = −30 5y − 5x = 10 4. 5y + 5x = −30 5y + 5x = 10

OpenStudy (astrophysics):

One problem at a time.

OpenStudy (anonymous):

those are the answer choices

OpenStudy (astrophysics):

I don't understand, what exactly is the question

OpenStudy (anonymous):

To rewrite y= 1/6x-5 and y=1/5x+2 as a system of equations

OpenStudy (anonymous):

I got A :)

OpenStudy (anonymous):

ohh

OpenStudy (astrophysics):

It's already a system of equation, but I see you just want to simplify it.

OpenStudy (astrophysics):

So you have \[y=\frac{ 1 }{ 6 }x-5\] so setting it up as they've showed you, so first move 1/6x to the left side how do you do that

OpenStudy (anonymous):

1/6x=1/6x-5 ?

OpenStudy (astrophysics):

?

OpenStudy (anonymous):

did you see the answer choices?

OpenStudy (astrophysics):

Yes

OpenStudy (anonymous):

was A right?

OpenStudy (astrophysics):

So for your first one you have \[y-\frac{ 1 }{ 6 }x=-5\]

OpenStudy (astrophysics):

Honestly, if you're not interested in the process tell me now.

OpenStudy (anonymous):

No, not really :) Sorry

OpenStudy (astrophysics):

Ok then, good luck!